# Unitary quantum gravitational physics and the CMB parity asymmetry

### v1

@@ Abstract
where $r=\big{|}\frac{1}{H}\big{|}\frac{1+UV}{1-UV}$ and $\frac{V}{U}=-e^{2Ht_{s}}$. The relation between the flat FLRW (4) and (8) is just a simple coordinate change Lanczos and Hoenselaers (1997); Hartman (2017). A very practiced view of flat dS spacetime (4) is that it covers only half of the dS space. But we must keep in mind the symmetry (6) which implies that (4) can cover the entire dS space. Thus the maximally symmetric nature of dS spacetime sets no difference whether it is expressed in flat, closed, open FLRW coordinates or the static coordinates Mukhanov and Winitzki (2007). The use of mapping flat coordinates to static coordinates here is that we can realize how unitarity can be achieved with DQFT whereas it is lost in the standard QFT as it was shown by Gibbons and Hawking in 1977 Gibbons and Hawking (1977). In Fig. 3, we show the unitary quantum physics in the static dS space through the conformal (not) Penrose diagram, where we see the three-dimensional space split by parity for the interior (the white region) and exterior (the dark region) of the cosmological event horizon. We can picture here how none of the observers lose any information beyond the horizon. If any state leaves the horizon it is only the direct-sum component of what is inside the horizon. This means any observer can perfectly reconstruct information about what is beyond the horizon. Any state beyond the horizon is not entangled with what is inside. This means the Hilbert spaces of the interior and exterior states cannot be connected by direct-product ($\otimes$) but rather should be related by direct-sum ($\oplus$). This has to be like this because beyond-the-horizon time is not the same, so one cannot do the same quantum mechanics everywhere. In other words, the gravitational horizons are special, they separate different spacetimes by a boundary. Thus, they cannot be treated as a standard null surface where quantum states can be entangled across. Furthermore, we stress that no entanglement across the horizon does not lead to firewalls which happens in standard QFTCS but not in direct-sum QFTCS. We suggest Ref. Sravan Kumar and Marto (2023b) to the reader for further analogous discussion on unitarity and no information loss in the context of the Schwarzschild black hole. In a nutshell, DQFT offers a new understanding of dS space where any observer’s physics of the universe is complete, and no information is lost outside any observer’s Universe. This is exactly what Schrödinger envisioned a quantum theory to achieve in 1956 Erwin Schrodinger (1956). Furthermore, our achievement of DQFT echoes well with the recent classical understanding of dS through the Black Hole Universe (BHU) proposal Gaztanaga (2022); Gaztañaga (2023).
@@
Direct-sum Inflation (DSI) and CMB parity asymmetry - Inflationary quantum fluctuations are the initial seeds for large-scale structure formation. It is important to understand their quantum nature to derive consistent CMB predictions. Inflation is by definition a quasi-dS expansion Starobinsky (1980) and this means that $\mathcal{P}\mathcal{T}$ (6) is spontaneously broken by the presence of a non-perturbative scalar field in addition to the GR’s tensor degree of freedom. Our focus here is on single-field inflation. The metric and scalar field matter quantum fluctuations are generated during inflation. The quantity which is a gauge invariant combination of them is called the curvature perturbation ($\zeta$), and it is an effective (scalar) degree of freedom. It connects the quantum physics of inflation to CMB observations. The action for curvature perturbation at the linear order is
@@
In the framework of DSI, the canonical variable is promoted to field operator that is split into direct-sum of two components
@@
in which they evolve forward and backward in time at the parity conjugate points in physical space. In analogy with (6), the time reversal operation (quantum mechanically) in an expanding quasi-dS Universe is
@@
Figure 4: The top panels show the anti-symmetric, $\mathcal{T}^{-}(\hat{n})=\left[\mathcal{T}(\hat{n})-\mathcal{T}(-\hat{n})\right]/2$, component of the Planck 2018 CMB temperature map Akrami et al. (2020a) in North and South polar galactic caps. The visualization is such that the Galactic plane, shown in Fig. 5, remains obscured along the edges. Parity conjugate positions can be identified by horizontally flipping into the opposite pole, where both caps have the same information with opposing signs: $\mathcal{T}^{-}(-\hat{n})=-\mathcal{T}^{-}(\hat{n})$. The bottom panel presents the observed $\mathcal{T}=\mathcal{T}^{+}+\mathcal{T}^{-}$ maps, including both the symmetric $\mathcal{T}^{+}(\hat{n})=\left[\mathcal{T}(\hat{n})+\mathcal{T}(-\hat{n})\right]/2$ and anti-symmetric components. The evident resemblance between the top and bottom panels indicates the prevalence of the anti-symmetric component, in perfect agreement with DSI predictions (21).
@@
$\mathcal{T}\left(\hat{n}\right)=\sum a_{\ell m}Y_{\ell m}\left(\hat{n}\right)\quad;\quad C_{\ell}=\frac{1}{2\ell+1}|a_{\ell m}^{2}|$ (14)
the Fourier coefficients $a_{\ell m}$ are characterized by the angular power spectrum $C_{\ell}$. Because $\mathcal{T}$ (and $\mathcal{P}\mathcal{T})$ is broken by the slow-roll parameters (13), we expect the quantum field during inflation to evolve time asymmetrically at parity conjugate points in physical space. This leads to an angular power spectrum $C_{\ell}$ of the CMB exhibiting an excess of power in the odd multipoles compared to the even ones, as described by (21) in Appendix C.
In Fig. 4 we can see visually the parity asymmetry by the close resemblance between odd parity temperature maps and the original CMB maps from Planck data Akrami et al. (2020a). This reflects that the Planck temperature maps contain an additional pure odd parity component. The DSI framework exactly generates this additional odd symmetry as we can see in (21). It is worth noticing here that if we take the average of even and odd power spectra we recover the well-known near-scale invariant power spectrum of Standard Inflation (SI). The benefit of DSI is that there are no additional free parameters. The parity asymmetry just depends on the spectral index $n_{s}\approx 0.9634\pm 0.0048$ at the scale $k_{\ast}=0.05\,{\rm Mpc}^{-1}$ ($\ell\approx 200$ or $\theta\lesssim 1^{\circ}$) as measured by the Planck data Akrami et al. (2020b).
In Fig. 6 we present how the observational data prefers significantly DSI over SI. Observed quantities include the observed low quadrupolar amplitude $C_{2}=C_{\ell=2}$, the even-to-odd multipole ratio of angular power spectra:
@@
and the mean $Z^{1}=<Z(\hat{n})>$ of the parity map $Z(\hat{n})\equiv\mathcal{T}(\hat{n})\mathcal{T}(-\hat{n})$ displayed in Fig.5, where $-\hat{n}=(\pi-\theta,\pi+\varphi)$ represents the parity conjugate or antipode of the sky position $\hat{n}=(\theta,\varphi)$. The likelihood probability associated with DSI in these plots is up to 650 times larger than that of SI. DSI favors odd parity, leading to a lower $C_{2}$. This, in turn, diminishes the significance of other CMB anomalies Schwarz et al. (2016), such as the Hemispherical Power Asymmetry, the axis of Evil, and the anomalous 2-point correlation function (See Gaztañaga and Sravan Kumar (2024) for further details).
Figure 5: Parity map: $Z(\hat{n})=\mathcal{T}(\hat{n})\mathcal{T}(-\hat{n})$ for Fig. 4 with galactic mask. This is a $\mathcal{P}$ conjugate image of itself. Positive values (red) correspond to the symmetric (even parity) component of $\mathcal{T}$, while negative values (blue) give the antisymmetric (odd parity) component. There is a clear excess of negative values, indicating a preference for odd parity. This together with Fig. 4 serves as a visual testament to the unitary QFTCS in nature.
Figure 6: Parity measurements derived from different mask (component separation) CMB temperature maps $\mathcal{T}(\hat{n})$ (dashed lines) against data simulations (grey areas represent 1, 2, and 3 sigma contours), alongside predictions from standard inflationary (SI) models (in red) and Direct-Sum inflation (DSI) predictions derived from DQFT (in blue).
Conclusions- In this letter, we underscore the pivotal role of unitary Quantum Field Theory in Curved Spacetime (QFTCS) in elucidating the intricacies of early Universe physics. Contrary to prevailing notions, we contend that the issue of unitarity in QFTCS is not inherently tied to quantum gravity at Planck scales, but rather stems from our incomplete comprehension of quantum fields in curved spacetime.
With the $\mathcal{P}\mathcal{T}$-symmetric formulation of Schrödinger equation (1), the DQFT framework reinstates unitarity, which was previously thought to be lost in curved spacetime. Fig. 2-3 visually represent this concept, illustrating how a hypothetical classical observer consistently experiences pure states without encountering information loss beyond the horizon.
Furthermore, the striking correlation illustrated in Fig. 4-6 between the observed hot and cold mirrored structures of the CMB map at antipodal points strongly supports the unitary treatment of inflationary quantum fluctuations within the framework of DQFT. Our findings indicate that superhorizon fluctuations at antipodes manifest as $\mathcal{P}\mathcal{T}$ conjugates, a prediction derived from direct-sum QFTCS. This prediction is substantially favored with a probability over 650 times greater than that of the standard QTFCS.
@@ Acknowledgements.
Appendix A Appendix A: Klein-Gordon field operator in DQFT
@@
which defines the positive norm states in the direct-sum vacuums $|0_{T}\rangle=|0_{+}\rangle\oplus|0_{-}\rangle,\,a_{(\pm)}|0_{\pm}\rangle=0$. The DQFT splits the Fock space as the direct-sum of two Fock spaces corresponding to the field’s components $\mathcal{F}_{T}=\mathcal{F}_{+}\oplus\mathcal{F}_{-}$ where states of forward in time and backward in time at parity conjugate regions of position space can be created and annihilated by the pair of operators (18). Since the Minkowski spacetime is $\mathcal{P}\mathcal{T}$ symmetric the scalar field operator in DQFT respects the same symmetry. For further details on DQFT and $\mathcal{C}\mathcal{P}\mathcal{T}$ invariance see Sravan Kumar and Marto (2023a).
@@ Appendix B Appendix B: DQFT in dS
Applying DQFT in dS Gaztañaga and Sravan Kumar (2024); Sravan Kumar and Marto (2023a), by construction the scalar field operator is the direct-sum of two components
@@
Here the operators $d_{\pm\,\textbf{k}},\,d^{\dagger}_{\pm\,\textbf{k}}$ satisfy the similar relations like in (18). The dS mode functions $\phi_{\pm\,k}=\frac{1}{\sqrt{2k}}\left(1\mp\frac{i}{k\tau}\right)e^{\mp ik\tau}$ corresponding to the standard Bunch-Davies vacuum of the DQFT.
Appendix C Appendix C: DSI angular power spectrum
Following Gaztañaga and Sravan Kumar (2024), the CMB power spectrum $C_{\ell}$ in DSI is:

### v2

@@ III DQFT in de Sitter & Unitarity ($G=\hbar=c=1$)
where $r=\big{|}\frac{1}{H}\big{|}\frac{1+UV}{1-UV}$ and $\frac{V}{U}=-e^{2Ht_{s}}$. The relation between the flat FLRW Eq.4 and Eq.12 is just a simple coordinate change (Lanczos and Hoenselaers (1997); Hartman (2017)). A very practiced view of flat dS spacetime in Eq.4 is that it covers only half of the dS space. However, we must keep in mind the symmetry in Eq.6, which implies that Eq.4 can cover the entire dS space. Thus, the maximally symmetric nature of the dS spacetime sets no difference whether it is expressed in flat, closed, open FLRW coordinates or the static coordinates (Mukhanov and Winitzki (2007)). The use of mapping flat coordinates to static coordinates here is so that we can realize how unitarity can be achieved with DQFT, whereas it is lost in standard QFT, as was shown by Gibbons and Hawking in 1977 Gibbons and Hawking (1977). In Fig. 2, we show unitary quantum physics in the static dS space through the (quantum) conformal diagram, where we see the three-dimensional space divided by parity for the interior and exterior of the cosmological event horizon. We can picture here how none of the observers loses any information beyond the horizon. If any state leaves the horizon, it is only the (complimentary) direct-sum component of what is inside the horizon. This means that any observer can perfectly reconstruct the information about what is beyond the horizon. According to DQFT, every density matrix of a maximally entangled pure state is split into direct-sum of 4 components
$\rho=\left(\rho_{I}\oplus\rho_{II}\right)\oplus\left(\rho_{III}\oplus\rho_{IV}\right)$ (13)
Each density matrix $\rho_{n},\,n=I,II,III,IV$ is defined by a pure state component in the corresponding (geometric) superselection sector Hilbert space $\mathcal{H}_{n}$. Since Von Neumann entropy of each component of density matrix vanishes $S_{n}=-Tr\rho_{n}\ln\rho_{n}=0$, any observer in regions I, II, III, IV (see Fig. 2) would witness pure states evolving into pure states. Thus, with DQFT we can have both unitarity and observer complementarity in dS (Kumar and Marto (2025)).
Figure 2: This picture depicts the (quantum) conformal diagram of static dS spacetime Eq.12 with four regions that are related by discrete transformations and the lines with arrows indicate the arrows of time in each region. Here a quantum field operator is written as direct-sum of four components $\hat{\phi}=\frac{1}{\sqrt{2}}\left(\hat{\phi}_{I}\oplus\hat{\phi}_{II}\right)\oplus\frac{1}{\sqrt{2}}\left(\hat{\phi}_{III}\oplus\hat{\phi}_{IV}\right)$ corresponding to the regions I to IV. This is not a Penrose diagram because each point represents half of the $S_{2}$ sphere in contrast to the usual Penrose diagram. This diagram is to represent quantum physics in dS space rather than classical physics for which the usual Penrose diagram is used.
According to DQFT, any state is spread across the horizon by distinct components in geometric superselection sectors. This implies that any information of the entangled state is shared by the sectorial Hilbert space, leading to pure states evolving into pure states. To be precise, the Hilbert spaces of the interior and exterior states cannot be connected by direct product ($\otimes$) but rather should be related by direct-sum ($\oplus$). This has to be like this because beyond-the-horizon time is not the same, so one cannot do the same quantum mechanics everywhere. In other words, the gravitational horizons are special, they separate different spacetimes by a boundary. Thus, they cannot be treated as a standard null surface, where quantum states can be entangled across. Furthermore, we stress that no entanglement across the horizon does not lead to firewalls, which happens in standard QFTCS but not in direct-sum QFTCS. For further analogous discussion on unitarity and no information loss in the context of the Schwarzschild black hole, see Kumar and Marto (2024). In a nutshell, DQFT offers a new understanding of dS space where any observer’s physics of the universe is complete, and no information is lost outside any observer’s universe. This is exactly what Schrödinger envisioned in 1956 for a quantum theory to achieve Erwin Schrodinger (1956). Furthermore, our achievement of DQFT echoes well with the recent classical understanding of dS through the Black Hole Universe (BHU) proposal Gaztanaga (2022); Gaztañaga (2023); Gaztanaga:2025cun.
IV CMB parity asymmetry
@@
Inflationary quantum fluctuations are the initial seeds for large-scale structure formation. It is important to understand their quantum nature to derive consistent CMB predictions. Inflation is by definition a quasi-dS expansion (Starobinsky (1980)) and this means that $\mathcal{P}\mathcal{T}$ in Eq.6 is spontaneously broken by the presence of a non-perturbative scalar field in addition to the GR’s tensor degree of freedom. Our focus here is on single-field inflation. The metric and scalar field matter quantum fluctuations are generated during inflation. The quantity which is a gauge-invariant combination of them is called the curvature perturbation ($\zeta$), and it is an effective (scalar) degree of freedom. It connects the quantum physics of inflation to CMB observations. The action for curvature perturbation at the linear order is
@@
In the framework of Direct-sum Inflation (DSI), the canonical variable is promoted to field operator that is split into direct-sum of two components
@@
in which they evolve forward and backward in time at the parity conjugate points in physical space. In analogy with Eq.6, the time reversal operation (quantum mechanically) in an expanding quasi-dS Universe is
@@
These sign differences in the slow-roll parameters, which break the time-reversal symmetry of de Sitter spacetime, lead to corrections in the evolution of quantum fluctuations $\hat{v}_{(\pm)}$ in the parity-conjugate regions. This results in an asymmetry in the CMB temperature under spatial parity.
In spherical coordinates, the parity transformation maps a point at radial distance $r$ from angular position $(\theta,\,\varphi)$ to its antipode, $(\pi-\theta,\,\pi+\varphi)$. This transformation cannot be realized by any rotation, and thus represents a discrete global symmetry distinct from statistical isotropy. It is important to emphasize that this global spatial parity asymmetry is unrelated to the observed birefringence in the CMB polarization, which arises from parity-violating interactions with well-defined transformation properties Komatsu:2022nvu.
Figure 3: The top panels show the anti-symmetric, $\mathcal{T}^{-}(\hat{n})=\left[\mathcal{T}(\hat{n})-\mathcal{T}(-\hat{n})\right]/2$, component of the Planck 2018 CMB temperature map (Akrami et al. (2020a)) in North and South polar galactic caps. The visualization is such that the Galactic plane, shown in Fig. 4, remains obscured along the edges. Parity conjugate positions can be identified by horizontally flipping into the opposite pole, where both caps have the same information with opposing signs: $\mathcal{T}^{-}(-\hat{n})=-\mathcal{T}^{-}(\hat{n})$. The bottom panel presents the observed $\mathcal{T}=\mathcal{T}^{+}+\mathcal{T}^{-}$ maps, including both the symmetric $\mathcal{T}^{+}(\hat{n})=\left[\mathcal{T}(\hat{n})+\mathcal{T}(-\hat{n})\right]/2$ and anti-symmetric components. The evident resemblance between the top and bottom panels indicates the prevalence of the anti-symmetric component, in perfect agreement with DSI predictions Eq.27.
@@
$\mathcal{T}\left(\hat{n}\right)=\sum a_{\ell m}Y_{\ell m}\left(\hat{n}\right)\quad;\quad C_{\ell}=\sum_{m}\frac{|a_{\ell m}|^{2}}{2\ell+1}$ (19)
the Fourier coefficients $a_{\ell m}$ are characterized by the angular power spectrum $C_{\ell}$. Because $\mathcal{T}$ (and $\mathcal{P}\mathcal{T})$ is broken by the slow-roll parameters in Eq.18, we expect the quantum field during inflation to evolve time asymmetrically at parity conjugate points in physical space. This leads to an angular power spectrum $C_{\ell}$ of the CMB exhibiting an excess of power in the odd multipoles compared to the even ones, as described by Eq.27 in Appendix C.
In Fig. 3, we can visually see the parity asymmetry by the close resemblance between odd-parity temperature maps and the original CMB maps from Planck data (Akrami et al. (2020a)). This reflects that the CMB temperature maps contain an additional pure odd-parity component. The DSI framework exactly generates this additional odd symmetry, as we can see in Eq.27. It is worth noticing here that if we take the average of even and odd power spectra, we recover the well-known near-scale invariant power spectrum of Standard Inflation (SI). The benefit of DSI is that there are no additional free parameters. The parity asymmetry only depends on the spectral index $n_{s}\approx 0.9634\pm 0.0048$ on the pivot scale $k_{\ast}=0.05\,{\rm Mpc}^{-1}$ ($\ell\approx 800$) as measured by Planck data (Akrami et al. (2020b)).
In Fig. 5 we present how the observational data significantly prefer DSI over SI. Observed quantities include the observed low quadrupolar amplitude $C_{2}=C_{\ell=2}$, the even-to-odd multipole ratio of angular power spectra:
@@
and the mean $Z^{1}=<Z(\hat{n})>$ of the parity map $Z(\hat{n})\equiv\mathcal{T}(\hat{n})\mathcal{T}(-\hat{n})$ displayed in Fig. 4, where $-\hat{n}=(\pi-\theta,\pi+\varphi)$ represents the parity conjugate or antipode of the sky position $\hat{n}=(\theta,\varphi)$. The likelihood probability associated with DSI in these plots is up to 650 times larger than that of SI. DSI favors odd parity, leading to a lower $C_{2}$. This, in turn, diminishes the significance of other CMB anomalies (Schwarz et al. (2016)), such as the Hemispherical Power Asymmetry, the axis of Evil, and the anomalous 2-point correlation function (see Gaztañaga and Kumar (2024) for further details).
Figure 4: Parity map: $Z(\hat{n})=\mathcal{T}(\hat{n})\mathcal{T}(-\hat{n})$ for Fig. 3 with galactic mask. This is a $\mathcal{P}$ conjugate image of itself. Positive values (red) correspond to the symmetric (even parity) component of $\mathcal{T}$, while negative values (blue) give the antisymmetric (odd parity) component. There is a clear excess of negative values, indicating a preference for odd parity. This, together with Fig. 3 serves as a visual testament to the unitary QFTCS in nature.
Figure 5: Parity measurements derived from different mask (component separation) CMB temperature maps $\mathcal{T}(\hat{n})$ (dashed lines) against data simulations (grey areas represent 1, 2, and 3 sigma contours), alongside predictions from standard inflationary (SI) models (in red) and Direct-Sum inflation (DSI) (in blue). The later is 650 more probable.
V Conclusions
In this paper, we underscore the pivotal role of unitary Quantum Field Theory in Curved Spacetime (QFTCS) in elucidating the intricacies of early Universe physics. Contrary to prevailing notions, we contend that the issue of unitarity in QFTCS is not inherently tied to quantum gravity at Planck scales, but rather stems from throwing away the $\mathcal{P}\mathcal{T}$ symmetry for quantum fields in curved spacetime.
With the $\mathcal{P}\mathcal{T}$-symmetric formulation of Schrödinger equation Eq.1, the DQFT framework reinstates unitarity, which was previously thought to be lost in curved spacetime. Fig. 1-2 visually represent this concept, illustrating how a hypothetical classical observer consistently experiences pure states without encountering information loss beyond the horizon. Furthermore, the striking correlation illustrated in Fig. 3-5 between the observed hot and cold mirrored structures of the CMB map at antipodal points strongly supports the unitary treatment of inflationary quantum fluctuations within the framework of DQFT. Our findings indicate that superhorizon fluctuations at antipodes manifest as $\mathcal{P}\mathcal{T}$ conjugates, a prediction derived from direct-sum QFTCS. This prediction is substantially favored with a probability over 650 times greater than that of the standard QTFCS.
@@ Acknowledgements.
Appendix A Klein-Gordon field operator in DQFT
@@
$[\hat{a}_{(\pm)\,\textbf{k}},\,\hat{a}^{\dagger}_{(\mp)\,\textbf{k}^{\prime}}]$ $=0$
which defines the positive norm states in the direct-sum vacuums $|0_{T}\rangle=|0_{+}\rangle\oplus|0_{-}\rangle=\begin{pmatrix}|0_{+}\rangle\\ |0_{-}\rangle\end{pmatrix},\,a_{(\pm)}|0_{\pm}\rangle=0$. The second condition in Eq.23 implies commutativity of operators $\left[\hat{\phi}_{+},\,\hat{\phi}_{-}\right]=0$, which can be seen as a new causality condition. The The Fock space in DQFT is the direct-sum of two Fock spaces (geometric superselection sectors) $\mathcal{F}_{T}=\mathcal{F}_{+}\oplus\mathcal{F}_{-}$ corresponding to the components of field operators ϕ^±subscript^italic-$\hat{\phi}_{\pm}$ that lead to the quantum states forward and backward in time at parity conjugate regions of position space. These field operators are defined through the pairs of creation and annihilation operators Eq.23. Since the Minkowski spacetime is $\mathcal{P}\mathcal{T}$ symmetric, the scalar field operator in DQFT respects the same symmetry. Generalizing this structure to all quantum fields in Minkowski spacetime, such as complex scalar, vector, and fermionic degrees of freedom, is very straightforward. Any standard model (single particle) state, whether it is particle ($|\rm SM\rangle)$ or anti-particle ($|\overline{\rm SM}\rangle$, is expressed in DQFT as direct-sum of two components
$|\rm SM\rangle=\begin{pmatrix}|{\rm SM}_{+}\rangle\\ |{\rm SM}_{-}\rangle\end{pmatrix},\quad|\overline{{\rm SM}}\rangle=\begin{pmatrix}|\overline{{\rm SM}}_{+}\rangle\\ |\overline{{\rm SM}}_{-}\rangle\end{pmatrix}$ (24)
with respect to $\mathcal{P}\mathcal{T}$ conjugate splitting of the vacuum $|0_{M}\rangle=|0_{M+}\rangle\oplus|0_{M-}\rangle$. For further details on DQFT in particular, the CPT invariance of scattering amplitudes which holds in both geometric superselection sectors, and quantizing other fields of the standard model, see Kumar and Marto (2025).
@@ Appendix B Appendix B: DQFT in dS
Applying DQFT in dS (Gaztañaga and Kumar (2024); Kumar and Marto (2025)), the canonically rescaled scalar field ($\phi\to e^{Ht}\phi$) operator is
@@
Here the operators $d_{\pm\,\textbf{k}},\,d^{\dagger}_{\pm\,\textbf{k}}$ satisfy the similar relations like in Eq.23. The dS mode functions $\phi_{\pm\,k}=\frac{1}{\sqrt{2k}}\left(1\mp\frac{i}{k\tau}\right)e^{\mp ik\tau}$ corresponding to the Bunch-Davies vacuum of the DQFT, which recover short distance behavior of quantum fields in Minkowski Eq.22.
Appendix C DSI angular power spectrum
Following Gaztañaga and Kumar (2024), the CMB power spectrum $C_{\ell}$ in DSI is: